OpenStudy (anonymous):

The length of a rectangle is four times its width. If the perimeter of the rectangle is 70yd , find its area.

OpenStudy (anonymous):

rectangle formulas used area perimeter

OpenStudy (anonymous):

this question is the same as the first one 2w+2(L+4)=70 2w+4=35 w=15.5yd L=19.5 area=15.5 x 19.5= 302.25 yd^2

OpenStudy (anonymous):

@LilySwan your answer is wrong

OpenStudy (anonymous):

the answer was 196 yds squared.

OpenStudy (anonymous):

how we solved the first one with the same formula you may try

OpenStudy (anonymous):

P= 2L +2W length is 4 x width L = 4 w W = W P = 70 with this information you can begin solving all you just need to do is plug in the information you should get and answer followed then you can use the area formula after finding length and width A=LW

OpenStudy (anonymous):

THE ANSWER WAS 196!

OpenStudy (anonymous):

im just telling you how to do this

OpenStudy (anonymous):

because if you verify it by finding the perimeter again you will know that the found length and width are correct